OpenStudy (anonymous):

how to find solve x^2 + y^2 = 25 with y= x-1

OpenStudy (gorv):

we have y=x-1 put this value of y in other equation

OpenStudy (gorv):

x^2+(x-1)^2=25

OpenStudy (gorv):

x^2+x^2+1-2x=25